OpenStudy (ashley1nonly):

f(x)=x^2 between x=2 and x=4 using two rectangles to estimate the area under f(x)= using four rectangles to estimate the area under f(x)=

OpenStudy (ashley1nonly):

15 and 15/2 are the answers i got

OpenStudy (campbell_st):

are you using upper rectangles or lower rectangles ..?

OpenStudy (ashley1nonly):

i thing lower

OpenStudy (ashley1nonly):

since it says under

OpenStudy (campbell_st):

ok... so it's like

OpenStudy (ashley1nonly):

delta x [(f x0)+(fx1) 4-2 / 2 =1 1(f(2)+f(3)) 1(4+9) =15 for the first one

OpenStudy (campbell_st):

slight error 4 + 9 = 13

OpenStudy (ashley1nonly):

for the one with 4 triangle 4-2 / 4 = 1/2 1/2(f(2)+f(3) 1/2 (13) 13/2

OpenStudy (campbell_st):

well with 4 rectangles each will have a width of 1/2  so you will have areas 1/2 x 2 = 1/2 x 25/4 = 1/2 x 3 = 1/2 x 49/4 = then add them all